# SCKansformer: Fine-Grained Classification of Bone Marrow Cells via Kansformer Backbone and Hierarchical Attention Mechanisms

### v1

@@ 1 Introduction
In recent years, the morbidity and mortality rates of malignant tumors, such as acute leukemia, have shown a clear upward trend worldwide [19]. In the diagnostic process, hospitals initially employ cytological examinations of peripheral blood smears and bone marrow smears to determine the specific type of malignant tumor. For instance, based on cytology examination results, doctors can identify the type of leukemia and formulate targeted treatment plans. In cytological examinations, medical personnel focus on the number and proportion of various types of white blood cells in a sample from a leukemia patient compared to a healthy individual, relying on the percentage of primitive cells to determine the type and progression of the disease. Thus, the accuracy of bone marrow or peripheral blood smear tests hinges on the ability to accurately count the different types of blood cells in the sample. The results of bone marrow smear cytology are especially critical for clinicians as they serve as essential references for diagnosing and treating the disease.
Within a healthy body, the number of white blood cells remains within a specific range [18]. However, the number and proportion of white blood cells can fluctuate in response to inflammation or blood disorders. Therefore, routine blood tests are a vital tool for assessing overall health and assisting doctors in diagnosis [20]. For example, a sudden increase in lymphocytes may indicate conditions such as whooping cough, a bacterial infection, or cancer of the lymphatic system. Although initial blood tests can provide direction for diagnosing a disease by evaluating the number and percentage of different white blood cells, they cannot definitively diagnose a disease. Consequently, peripheral blood smears are often necessary for accurate diagnosis and treatment. This cost-effective diagnostic method is generally employed when routine blood tests yield abnormal results and is designed to detect morphological changes in all types of blood cells in peripheral blood. By examining the morphology and number of white blood cells, red blood cells, and platelets, one can gain a better understanding of the underlying cause of a disease.
@@
With the rapid advancement of computer hardware, deep learning has made significant strides in computer vision [21]. Increasingly, researchers are adopting deep learning models to study blood cell classification. Compared to traditional machine learning methods, deep learning models simplify the processes of image processing and feature extraction, making them more automated. However, cell examination in bone marrow blood microscopy images remains a complex challenge [22]. The accurate classification of multiple similar classes of blood cells is particularly difficult due to the large number of cell types and the presence of similar features among different cell classes.
@@
To address these challenges of automatic classification and recognition of blood cells in bone marrow microscopic images, this paper proposes a fine-grained classification model for bone marrow blood cells, SCKansformer. This approach utilizes the Kolmogorov-Arnold network to significantly enhance the model’s nonlinear feature expression capabilities and interpretability. By optimizing feature representation through spatial-channel reconstruction convolution, the method reduces feature redundancy, making the model more efficient in handling high-dimensional microscopic image data. Additionally, the integration of a global-local attention encoding mechanism effectively captures both the explicit and implicit features of bone marrow blood cells, enabling accurate classification of nearly forty types of bone marrow blood cells. This method not only provides essential reference standards for clinicians to diagnose and treat major dangerous diseases but also further enhances clinicians’ ability to diagnose acute leukemia and other significant hematological diseases, thus holding substantial significance in clinical auxiliary diagnosis. The primary contributions of this paper are as follows:
@@ 2.2 Machine Learning Blood Cell Classification Methods
In the 1980s, numerous researchers, both domestically and internationally, began exploring the use of computer image processing techniques and traditional machine learning algorithms to simulate the analysis of leukemia cell morphology under a microscope. Researchers employed various methods to process images of white blood cells. For instance, Ramoser et al. [1] initially applied K-means clustering to identify the nucleus location. They then used the MSER algorithm to threshold the image at different levels, followed by further differentiation of the nucleus and cytoplasm using K-means clustering. Based on the segmented leukocytes, they extracted color and shape features and subsequently classified the leukocytes using a Support Vector Machine (SVM). Similarly, Duan et al. [2] utilized the SMACC algorithm to extract leukocyte regions and morphological operations to determine leukocyte locations. They employed the ISODATA clustering algorithm for further segmentation to identify the nucleus region. Based on this segmentation, they extracted spectral, texture, and shape features and classified the leukocytes using an SVM.
@@
These early attempts at automating leukocyte classification highlighted the potential for computer-assisted analysis but also underscored the need for continued advancements in feature extraction, dimensionality reduction, and classification techniques.
@@ 3.1 Overall Architecture
The structure of the SCKansformer model is illustrated in Fig. 1. The model comprises three main components: the Kansformer Encoder, the SCConv Encoder, and the Global-Local Attention Encoder (GLAE). The Kansformer Encoder enhances the efficiency and interpretability of the Vision Transformer (ViT) model [25] by replacing the traditional MLP layer [36, 37, 38] with a Kolmogorov-Arnold Network (KAN) [16]. This substitution improves the nonlinear feature representation and interpretability of the ViT model, with the structure consisting of alternating Multi-head Self-Attention (MSA) layers and KAN networks. The SCConv Encoder comprises a Spatial Reconstruction Unit (SRU) and a Channel Reconstruction Unit (CRU). These units are designed to reduce redundant information in the features extracted from CNNs, specifically targeting spatial and channel redundancy through separation-reconstruction and separation-transformation fusion strategies, respectively. The GLAE consists of an MSA module and a Local Part module for capturing long-range dependencies and local features. The MSA module learns long-range dependencies through the self-attention mechanism, while the Local Part module extracts local features between neighboring pixels using depth-separable convolution. Through the synergistic work of these three components, the SCKansformer model effectively processes microscopic images, thereby enhancing both the performance and interpretability of the model. This enables the accurate classification of bone marrow blood cell images.
@@ 3.2 Kansformer Encoder
$\bm{\Phi}=\left\{\phi_{q,p}\right\},\quad p=1,2,\cdots,n_{\text{in }},\quad q=1,2\cdots,n_{\text{out }}.$ (3)
The computation result of the KAN network from layer $k$ to layer $k+1$ can be expressed in matrix form, as shown in Equation (4):
$\mathbf{Z}_{k+1}=\underbrace{\left(\begin{array}[]{cccc}\phi_{k,1,1}(\cdot)&\phi_{k,1,2}(\cdot)&\cdots&\phi_{k,1,n_{k}}(\cdot)\\ \phi_{k,2,1}(\cdot)&\phi_{k,2,2}(\cdot)&\cdots&\phi_{k,2,n_{k}}(\cdot)\\ \vdots&\vdots&&\vdots\\ \phi_{k,n_{k+1},1}(\cdot)&\phi_{k,n_{k+1},2}(\cdot)&\cdots&\phi_{k,n_{k+1},n_{k}}(\cdot)\end{array}\right)}_{\bm{\Phi}_{k}}\mathbf{Z}_{k}.$ (4)
@@ 3.3 Global-Local Attention Encoder
In addition, variations in color and quality of microscopic images often arise due to differences in equipment and imaging conditions. Furthermore, the long-tailed distribution and low resolution inherent in the myeloid blood cell dataset pose significant challenges, limiting the model’s ability to recognize and classify rare cell classes. As depicted in Fig. 1, the GLAE primarily consists of a MSA module and a Local Part module. The GLAE is designed to address these challenges by focusing not only on the long-range dependencies within microcellular images but also by effectively extracting local features through depth-separable convolution. This dual approach allows the model to capture complex relationships between neighboring pixels, enhancing its ability to discern subtle differences in cell morphology.
@@ 3.3.1 MSA Module
where the weight matrices $W_{Q},W_{K},$ and $W_{V}$ are elements of $\mathbb{R}^{d\times d}$ and $Q$ denotes the query information, $K$ denotes the key information (used to match $Q$), and $V$ represents the content vector. As shown in Equations (5) and (6), the inner product of $Q$ and $K$ (i.e., $QK^{T}$) captures the interrelated relationships between feature points. This product is then scaled by $\sqrt{d}$ where $d$ is the dimension of the hidden layer. This scaling serves two purposes:
1). To prevent the values from becoming too large when subsequently using the Softmax function, which could result in partial derivatives tending toward zero. 2). To normalize $QK^{T}$, ensuring it has an expected distribution with a mean of 0 and a variance of 1.
@@ 3.4 SCConv Encoder
Figure 3: The framework of of SCConv Encoder. The SCConv Encoder optimizes the feature representation by means of Spatial Reconstruction Units and Channel Reconstruction Units, which reduces feature redundant information of microscopic images.
@@ 3.4.2 CRU Module
Experiment
@@ 3.5 Data Acquisition and Processing
Due to the difficulties associated with collecting and labeling images of bone marrow blood samples, and the highly granular classification required for bone marrow blood cells, there is a notable lack of accurately labeled, large-scale, fine-grained bone marrow blood cell datasets in the academic community. To address this gap, our team, in collaboration with the Department of Hematology at Zhejiang Hospital in Hangzhou, Zhejiang Province, China, has established the private BMCD-FGCD dataset. This dataset, which contains over 10,000 images, comprises nearly 40 different classifications and has been accurately labeled by three senior physicians with more than 20 years of experience in hematology.
The BMCD-FGCD dataset includes a total of 92,335 bone marrow blood cell images that span nearly 40 distinct blood cell categories. For the purposes of this study, the dataset is divided into a training set containing 73,877 samples and a test set containing 18,458 samples. Detailed statistical information is presented in Fig. 4. In an effort to contribute positively to the advancement of the field, we have decided to make the BMCD-FGCD dataset publicly available to other researchers.
@@
The BMCD-FGCD dataset comprises original images derived from bone marrow sample images within the hospital data system. These images have been collected from patients who presented to the Department of Hematology at Zhejiang Hospital and were deemed by specialized physicians to require bone marrow aspiration for diagnostic assistance following preliminary clinical evaluations. The construction of the BMCD-FGCD dataset, as illustrated in Fig. 5, was methodically conducted through a tripartite process:
Figure 5: Workflow of the establishment of the proposed dataset. (a) Data were extracted from patients’ bone marrow fluid samples obtained at the Department of Hematology at Zhejiang Hospital. The preparation and examination of the smears were carried out using two types of Olympus microscopes. (b) Images deemed to be of low quality underwent standardization through image processing software and systems to ensure their suitability for algorithmic processing. (c) Expert physicians at Zhejiang Hospital meticulously identified and manually annotated each image with the respective cell type. Ultimately, nearly 40 types of bone marrow blood cells were classified within the dataset.
@@ 3.5.1 Acquisition Stage
Patients are positioned either prone or laterally under the careful supervision of physicians. The skin overlying the puncture site is thoroughly disinfected, and a local anesthetic is administered to achieve targeted infiltration. Utilizing a sterile, disposable puncture device equipped with a micropuncture needle, the clinician carefully advances the needle through the cortical bone and into the marrow cavity. Upon successful entry into the marrow cavity, the physician withdraws the needle’s core and attaches a 10 ml syringe to aspirate the bone marrow fluid, which is subsequently prepared as a smear sample.
After being dried and stained, the smear samples are examined cytologically under a microscope by a specialized physician. The cell images and their contents are then entered into Zhejiang Hospital’s electronic medical record system. Under the supervision and guidance of specialized physicians, our team selected high-quality bone marrow blood images of patients of different ages, genders, and diseases from Zhejiang Hospital’s electronic medical record system (image archiving system), ensuring the privacy of the data sources throughout the process.
@@ 3.6 Comparison to Existing Datasets
ALL-IDB dataset [31]: This dataset is a specialized resource developed for the segmentation and classification of mother cells (immature or precursor cells) in Acute Lymphocytic Leukemia (ALL). The images in the dataset are captured using a Canon PowerShot G5 camera, saved in JPG format with a 24-bit color depth, and the dataset is divided into two subsets: ALL-IDB1, which is dedicated to segmentation, and ALL-IDB2, which is tailored for classification tasks. ALL-IDB1 comprises 108 labeled peripheral blood smear images with a resolution of 2,592 × 1,944 pixels, featuring approximately 39,000 blood cells, including 510 lymphocytes from ALL patients validated by expert oncologists. ALL-IDB2 contains images cropped from the ALL-IDB1 collection that specifically display acute lymphoblasts, each standardized to a resolution of 257 × 257 pixels. This subset includes 260 images, of which half portray lymphoblasts from ALL patients.
@@ 3.7 Experimental Settings
For the training data, input images were first randomly cropped to a size of $224\times 224$ pixels, and data augmentation techniques, such as random horizontal flipping, were applied to enhance the dataset. The Adam optimizer was employed in conjunction with a cosine annealing algorithm to prevent the model from converging to local optima.
To verify the validity of the model during the validation process, images were first resized to $256\times 256$ pixels and then center-cropped to $224\times 224$ pixels. This preprocessing step ensured consistency and accuracy in the validation results.
@@ 3.8 Evaluation Indicators
To comprehensively evaluate the performance of different models on the category-unbalanced BMCD-FGCD dataset, we use the following standard evaluation metrics: Accuracy, Precision, Recall, and F1-score. These metrics are calculated using the formulas shown below:

### v2

@@ 1 Introduction
In recent years, the morbidity and mortality rates of malignant tumors, such as acute leukemia, have shown a clear upward trend worldwide [18]. In the diagnostic process, hospitals initially employ cytological examinations of peripheral blood smears and bone marrow smears to determine the specific type of malignant tumor. For instance, based on cytology examination results, doctors can identify the type of leukemia and formulate targeted treatment plans [37]. In cytological examinations, medical personnel focus on the number and proportion of various types of white blood cells in a sample from a leukemia patient compared to a healthy individual, relying on the percentage of primitive cells to determine the type and progression of the disease. Thus, the accuracy of bone marrow or peripheral blood smear tests hinges on the ability to accurately count the different types of blood cells in the sample. The results of bone marrow smear cytology are especially critical for clinicians as they serve as essential references for diagnosing and treating the disease.
Within a healthy body, the number of white blood cells remains within a specific range. However, the number and proportion of white blood cells can fluctuate in response to inflammation or blood disorders. Therefore, routine blood tests are a vital tool for assessing overall health and assisting doctors in diagnosis [19]. For example, a sudden increase in lymphocytes may indicate conditions such as whooping cough, a bacterial infection, or cancer of the lymphatic system. Although initial blood tests can provide direction for diagnosing a disease by evaluating the number and percentage of different white blood cells, they cannot definitively diagnose a disease. Consequently, peripheral blood smears are often necessary for accurate diagnosis and treatment. This cost-effective diagnostic method is generally employed when routine blood tests yield abnormal results and is designed to detect morphological changes in all types of blood cells in peripheral blood. By examining the morphology and number of white blood cells, red blood cells, and platelets, one can gain a better understanding of the underlying cause of a disease.
@@
With the rapid advancement of computer hardware, deep learning has made significant strides in computer vision. Increasingly, researchers are adopting deep learning models to study blood cell classification. Compared to traditional machine learning methods, deep learning models simplify the processes of image processing and feature extraction, making them more automated. However, cell examination in bone marrow blood microscopy images remains a complex challenge [20]. The accurate classification of multiple similar classes of blood cells is particularly difficult due to the large number of cell types and the presence of similar features among different cell classes.
@@
To address these challenges of automatic classification and recognition of blood cells in bone marrow microscopic images, this paper proposes a fine-grained classification model for bone marrow blood cells, SCKansformer. This method not only provides essential reference standards for clinicians to diagnose and treat major dangerous diseases but also further enhances clinicians’ ability to diagnose acute leukemia and other significant hematological diseases, thus holding substantial significance in clinical auxiliary diagnosis. The primary contributions of this paper are as follows:
@@ 2.2 Machine Learning Blood Cell Classification Methods
In the 1980s, numerous researchers began exploring the use of computer image processing techniques and traditional machine learning algorithms to simulate the analysis of leukemia cell morphology under a microscope. Researchers employed various methods to process images of white blood cells. For instance, Ramoser et al. [1] initially applied K-means clustering to identify the nucleus location. They then used the MSER algorithm to threshold the image at different levels, followed by further differentiation of the nucleus and cytoplasm using K-means clustering. Based on the segmented leukocytes, they extracted color and shape features and subsequently classified the leukocytes using a Support Vector Machine (SVM). Similarly, Duan et al. [2] utilized the SMACC algorithm to extract leukocyte regions and morphological operations to determine leukocyte locations. They employed the ISODATA clustering algorithm for further segmentation to identify the nucleus region. Based on this segmentation, they also tried to extract spectral, texture, and shape features and classified the leukocytes using an SVM.
@@ 3.1 Overall Architecture
The structure of the SCKansformer model is illustrated in Fig. 1. The model comprises three main components: the Kansformer Encoder, the SCConv Encoder, and the Global-Local Attention Encoder (GLAE). The Kansformer Encoder enhances the efficiency and interpretability of the Vision Transformer (ViT) model [23] by replacing the traditional MLP layer [34] with a Kolmogorov-Arnold Network (KAN) [16]. This substitution improves the nonlinear feature representation and interpretability of the ViT model, with the structure consisting of alternating Multi-head Self-Attention (MSA) layers and KAN networks. The SCConv Encoder comprises a Spatial Reconstruction Unit (SRU) and a Channel Reconstruction Unit (CRU). These units are designed to reduce redundant information in the features extracted from CNNs, specifically targeting spatial and channel redundancy through separation-reconstruction and separation-transformation fusion strategies, respectively. The GLAE consists of an MSA module and a Local Part module for capturing long-range dependencies and local features. The MSA module learns long-range dependencies through the self-attention mechanism, while the Local Part module extracts local features between neighboring pixels using depth-separable convolution. Through the synergistic work of these three components, the SCKansformer model effectively processes microscopic cell images, thereby enhancing both the performance and interpretability of the model. This makes accurate classification of bone marrow blood cell images a possibility.
@@ 3.2 Kansformer Encoder
$\bm{\Phi}=\left\{\phi_{k,q,p}\right\},\quad p=1,2,\cdots,n_{\text{in }},\quad q=1,2\cdots,n_{\text{out }}.$ (3)
The pre-activation of $\phi_{k,q,p}$ is simply $Z_{k,p}$; the post-activation of $\phi_{k,q,p}$ is denoted by $\tilde{Z}_{k,q,p}\equiv\phi_{k,q,p}(Z_{k,p})$. The activation value of the $(k+1,q)$ neuron is simply the sum of all incoming post-activations, as shown in Equation(4):
$Z_{k+1,q}=\sum_{p=1}^{n_{k}}\tilde{Z}_{k,q,p}=\sum_{p=1}^{n_{k}}\phi_{k,q,p}(Z_{k,p}),\qquad q=1,\cdots,n_{out}.$ (4)
@@ 3.3 Global-Local Attention Encoder
The variations in color and quality of microscopic images often arise due to differences in equipment and imaging conditions. Furthermore, the long-tailed distribution and low resolution inherent in the myeloid blood cell dataset pose significant challenges, limiting the model’s ability to recognize and classify rare cell classes. As depicted in Fig. 1, the core design of the GLAE integrates the MSA module to capture long-range dependencies and the Local Part module to extract local features from neighboring pixels. These components are harmonized through an adaptive weight-sharing mechanism, ensuring that both global and local features are processed seamlessly throughout the network. This balanced approach enables the model to address the dataset’s complexities by capturing extensive dependencies while effectively extracting detailed local features using depth-separable convolution. This dual approach allows the model to capture complex relationships between neighboring pixels, enhancing its ability to discern subtle differences in cell morphology.
@@ 3.3.1 MSA Module
where the weight matrices $W_{Q},W_{K},$ and $W_{V}$ are elements of $\mathbb{R}^{d\times d}$ and $Q$ denotes the query information, $K$ denotes the key information (used to match $Q$), and $V$ represents the content vector. As shown in Equations (5) and (6), the inner product of $Q$ and $K$ (i.e., $QK^{T}$) captures the interrelated relationships between feature points. This product is then scaled by $\sqrt{d}$ where $d$ is the dimension of the hidden layer. This scaling serves two purposes: 1). To prevent the values from becoming too large when subsequently using the Softmax function, which could result in partial derivatives tending toward zero. 2). To normalize $QK^{T}$, ensuring it has an expected distribution with a mean of 0 and a variance of 1.
@@ 3.4 SCConv Encoder
Figure 3: The framework of SCConv Encoder. The SCConv Encoder optimizes the feature representation by means of Spatial Reconstruction Units and Channel Reconstruction Units, which reduces feature redundant information of microscopic images.
@@ 3.4.2 CRU Module
The primary advantage of the SCConv Encoder is its ability to target specific spatial and channel redundancies in microscopic images for feature compression. Unlike traditional dimensionality reduction techniques that mainly rely on statistical variance for feature reduction, the SCConv Encoder excels in preserving the spatial structural information crucial for accurate classification. Traditional methods such as PCA and Autoencoders, while effective, often struggle to maintain this essential information. In contrast, the SCConv Encoder significantly enhances classification accuracy by effectively reducing redundant information, particularly in high-dimensional microscopic image data. These benefits are further illustrated through visualizations, as shown in Fig. 4.
Figure 4: The feature dimensionality reduction visualizations of SCConv Encoder and other traditional methods. (a) Heatmap of PCA. (b) Heatmap of Autoencoder. (c) Heatmap of SCConv Encoder.
4 Experiment
@@ 4.1 Data Acquisition and Processing
Due to the difficulties associated with collecting and labeling images of bone marrow blood samples, and the highly granular classification required for bone marrow blood cells, there is a notable lack of accurately labeled, large-scale, fine-grained bone marrow blood cell datasets in the academic community. To address this gap, our team, in collaboration with the Department of Hematology at Zhejiang Hospital in Hangzhou, Zhejiang Province, China, has established the private BMCD-FGCD dataset. This research study was conducted retrospectively and in accordance with the principles of the Declaration of Helsinki. We applied for an informed consent waiver. Approval was granted by the Ethics Review Committee of Zhejiang Hospital: 2024 Proc. No. (023K).
This dataset, which contains over 10,000 images, comprises nearly 40 different classifications and has been accurately labeled by three senior physicians with more than 20 years of experience in hematology. The BMCD-FGCD dataset includes a total of 92,335 bone marrow blood cell images that span nearly 40 distinct blood cell categories. For the purposes of this study, the dataset is divided into a training set containing 73,877 samples and a test set containing 18,458 samples. Detailed statistical information is presented in Fig. 5. In an effort to contribute positively to the advancement of the field, we have decided to make the BMCD-FGCD dataset publicly available to other researchers.
@@
Our private dataset comprises original images derived from bone marrow sample images within the hospital data system. These images have been collected from patients who presented to the Department of Hematology at Zhejiang Hospital and were deemed by specialized physicians to require bone marrow aspiration for diagnostic assistance following preliminary clinical evaluations. The construction of the BMCD-FGCD dataset, as illustrated in Fig. 6, was methodically conducted through a tripartite process:
Figure 6: Workflow of the establishment of the proposed dataset. (a) Data were extracted from patients’ bone marrow fluid samples obtained at the Department of Hematology at Zhejiang Hospital. The preparation and examination of the smears were carried out using two types of Olympus microscopes. (b) Images deemed to be of low quality underwent standardization through image processing software and systems to ensure their suitability for algorithmic processing. (c) Expert physicians at Zhejiang Hospital meticulously identified and manually annotated each image with the respective cell type.
@@ 4.1.1 Acquisition Stage
Patients are positioned either prone or laterally under the careful supervision of physicians. The skin overlying the puncture site is thoroughly disinfected, and a local anesthetic is administered to achieve targeted infiltration. Utilizing a sterile, disposable puncture device equipped with a micropuncture needle, the clinician advances the needle through the cortical bone and into the marrow cavity. Upon successful entry into the marrow cavity, the physician withdraws the needle’s core and attaches a 10 ml syringe to aspirate the bone marrow fluid, which is subsequently prepared as a smear sample. After being dried and stained, the smear samples are examined cytologically under a microscope by a specialized physician. We selected high-quality images of patients of different ages, genders, and diseases from Zhejiang Hospital’s electronic medical record system.
@@ 4.2 Comparison to Existing Datasets
ALL-IDB dataset [29]: This dataset is a specialized resource developed for the segmentation and classification of mother cells (immature or precursor cells) in Acute Lymphocytic Leukemia (ALL). The images in the dataset are captured using a Canon PowerShot G5 camera, saved in JPG format with a 24-bit color depth, and the dataset is divided into two subsets: ALL-IDB1, which is dedicated to segmentation, and ALL-IDB2, which is tailored for classification tasks. ALL-IDB1 comprises 108 labeled peripheral blood smear images with a resolution of 2,592 × 1,944 pixels, featuring approximately 39,000 blood cells, including 510 lymphocytes from ALL patients validated by expert oncologists. ALL-IDB2 contains images cropped from the ALL-IDB1 collection that display acute lymphoblasts, each standardized to a resolution of 257 × 257 pixels. This subset includes 260 images, of which half portray lymphoblasts from ALL patients.
@@ 4.3 Experimental Settings
For the training data, input images were randomly cropped to $224\times 224$ pixels, and data augmentation techniques, such as random horizontal flipping, were applied to enhance the dataset. The Adam optimizer was used in conjunction with a cosine annealing algorithm. At the same time, for the testing data, images were resized to $256\times 256$ pixels and then center-cropped to $224\times 224$ pixels.
@@ 4.4 Evaluation Indicators
To comprehensively evaluate the performance of different models on the category-unbalanced BMCD-FGCD dataset, we use the following standard evaluation metrics: Accuracy, Precision, Recall, F1-score, Balanced Accuracy, and MCC. These metrics are calculated using the formulas shown below:
@@
Balanced $\text{Balanced Accuracy}=\frac{1}{2}(\frac{TP}{TP+FN}+\frac{TN}{TN+FP}),$ (25)
$\text{MCC}=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}},$ (26)